OpenStudy (anonymous):

write the equation of the line in standard form containing (9,8) and perpendicular to 3x-4y-5=0

OpenStudy (stefrheart):

so first you should put the equation into standard form

OpenStudy (anonymous):

Using Ax+By=c

OpenStudy (anonymous):

?

OpenStudy (texaschic101):

put your equation in y = mx + b form because we need to find the slope(m)

OpenStudy (texaschic101):

can you do that ?

OpenStudy (anonymous):

Is it -4=3x+5?

OpenStudy (texaschic101):

3x - 4y - 5 = 0 (put in y = mx + b form) -- get the y term on one side and everything else on the other side -4y = -3x + 5 y = 3/4x - 5/4 are you still with me or have I lost you ?

OpenStudy (anonymous):

Okay so you just add 4 to both sides .

OpenStudy (texaschic101):

no.. let me start again and I will go step by step 3x - 4y - 5 = 0 --- subtract 3x from both sides 3x - 3x - 4y - 5 = 0 - 3x -- combine like terms -4y - 5 = -3x -- add 5 to both sides -4y - 5 + 5 = -3x + 5 -- combine like terms -4y = -3x + 5 -- divide both sides by -4 y = (-3/-4)x + 5/(-4) y = 3/4x - 5/4 understand ?

OpenStudy (anonymous):

I understand a little. you lost me at step 4....

OpenStudy (texaschic101):

now that we have it in y = mx + b form......we know that the slope is in the m position. So our slope of y = 3/4x - 5/4 is 3/4. But we are looking for a perpendicular line...that means we need the negative reciprocal slope. All that means is " flip " the slope and change the sign. So the new slope is -4/3 (see how I flipped the slope and changed the sign) now we use y = mx + b...we already know the slope, so we need to find the y intercept(b) y = mx + b slope(m) = -4/3 (9,8)..x = 9 and y = 8 now we sub 8 = -4/3(9) + b 8 = - 12 + b 8 + 12 = b 20 = b so your perpendicular equation is : y = -4/3x + 20

OpenStudy (texaschic101):

oh...sorry.. whatever you do on one side of the equal sign, you have to do on the other. 3x - 4y - 5 = 0 --- subtract 3x from both sides -4y - 5 = -3x -- add 5 to both sides -4y = -3x + 5 -- now divide by -4 on both sides y = 3/4x - 5/4 where am I losing you ?

OpenStudy (texaschic101):

oh wait...you need it in standard form Ax + By = C y = -4/3x + 20 -- multiply everything by 3 3y = -4x + 60 -- add 4x to both sides 4x + 3y = 60 <=== standard form

OpenStudy (anonymous):

Is the top info right for the problem or is the last statement you wrote is what I really need?

OpenStudy (texaschic101):

the top info is just doing the problem...but I didn't realize you needed it in standard form, so the last part was me finishing the problem. The answer you need is : 4x + 3y = 60 -- thats your perpendicular line going through points (9,8)

OpenStudy (anonymous):

Okay i have an under standing. Of what you just showed me . But do I really need to do those first steps all the time when I have problems like this? Mathematics is my weakest subject. Im just trying to get an clear understanding.

OpenStudy (texaschic101):

yes...unfortunately you do. You have to do the first steps to find the slope. Then, for a perpendicular line, you have to find the negative reciprocal slope. Then you have to do your subbing in the formula y = mx + b to find your line in slope intercept form. And then you can convert it to standard form.